OpenStudy (anonymous):

A car is traveling at 118 km/h. What is its speed in mm/h?

OpenStudy (aaronq):

convert km to mm 1 km = 1 000 000 mm

OpenStudy (anonymous):

Is there a way u could show me in the dimensional analysis?

OpenStudy (aaronq):

\(118 \;km*\dfrac{1,000,000 mm}{1\;km}\) then just add "per hour", \h, to the answer

OpenStudy (anonymous):

So its 118000000 right?

OpenStudy (aaronq):

don't forget the units

OpenStudy (anonymous):

Yea ok

OpenStudy (anonymous):

Is that it?

OpenStudy (aaronq):

you tell me, do you think that answers the question?

OpenStudy (anonymous):

Yes but wat abt mm/h?

OpenStudy (aaronq):

what about it?

OpenStudy (anonymous):

Can u please show me how to do it step by step in dimensional analysis while explaning?

OpenStudy (aaronq):

we already did it. We found how many mm are in said km, since the time units are the same, there is nothing else to be done.

OpenStudy (anonymous):

Oooh lol sorry. I was thinking of the example the teacher showed me today. Sorry

OpenStudy (aaronq):

it's okay. if you were also changing the time intervals (say hours to seconds) then you would have do further conversions.

OpenStudy (anonymous):

Can u show me an example?

OpenStudy (anonymous):

Please

OpenStudy (aaronq):

1 km/h to 1 m/sec \(1\; km/hr*\dfrac{1000\;\color{red}{m}}{1\;km}*\dfrac{1\;hr}{60\;min}*\dfrac{1\;min}{60\;\color{red}{sec}}\) notice the units are cancelling out, and you only end up with m/sec

OpenStudy (aaronq):

you're essentially cancelling out units by multiplying by its values. 60 min in 1 hr 1000 m in 1 km etc.

OpenStudy (anonymous):

Ooh k that makes sense. Got it thanx!! Btw do u kno an app for this?

OpenStudy (aaronq):

ok cool. no, i'm sure if you search there'll be one though.

OpenStudy (anonymous):

Alright...im moving to algebra 2 can u help me please?

OpenStudy (aaronq):

I gotta go in a few min, sorry. You should ask those questions in the math section, though.

OpenStudy (anonymous):

K thanxx!!!